OpenStudy (anonymous):

how do I write (2x10^3) (3x10^2) in scientific notation?

OpenStudy (anonymous):

Also please tell me how to do it, because I have no idea.

OpenStudy (anonymous):

2(1000) * 3 (100) = 6(100000) = 6*10^5

OpenStudy (anonymous):

It's just like multiplying normally, add the exponents and multiply the constants.

OpenStudy (anonymous):